28th May (Thursday)

  • Operations dominated by MC jobs and user analysis; Preparation for data receiving
  • T0
  • T1

21th May (Thursday)

  • Operations dominated by MC jobs and user analysis
  • LHCb COmputing workshop going on, so not much operation activity
  • Downtime today though due to the Oracle DB upgrade
  • T0:
  • T1
    • Problem to contact SARA SRM Seems to be back. They clame that the fetch-crl problem could be now on CERN side. Upgrading on FTS servers and our vobox planned.

18th May (Monday)

15th May (Friday)

  • Operations dominated by MC jobs and user analysis
  • T0:
  • T1
    • Problem to contact SARA SRM (GGUS:113324) fetch-crl 3.0.16 seems to have solved the issue. Ticket closed.

11th May (Monday)

  • Operations dominated by MC jobs and user analysis
  • T0:
  • T1
    • Problem to contact SARA SRM (GGUS:113324) Waiting for NLT1 to deploy fetch-crl 3.0.16 (only in EPEL-testing)

7th May (Thursday)

  • Operations dominated by MC jobs and user analysis
  • T0:
  • T1
    • Problem to contact SARA SRM (GGUS:113324) no news from SARA

4th May (Monday)

  • Operations dominated by MC jobs and user analysis
  • T0:
  • T1

-- JoelClosier - 2016-01-07

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2016-01-07 - JoelClosier
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LHCb All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback